If a regular polygon has (N-10) sides, where N= (40/10)^2, what is the measure if one of its angles?
A) 120
B) 160
C)240
D)720

Answer by greenestamps(13200)



N = (40/10)^2 = 4^2 = 16

N-10 = 16-10 = 6

What is the measure of a an angle of a regular hexagon?

Note answer choices C and D are impossible....